Two equally charged, 1.00 g spheres are placed with 2.00 cm between their centers. When released, each begins to accelerate at 225 m/s2.

What is the magnitude of the charge on each sphere?
Answer #1
Coulomb's Law Equation


F = (k * Q1 *Q2) / r^2 or ma = (k * Q1 *Q2) / r^2


k = 9 x 10^9 (Constant)


F = electrostatic force on release = 9*10^9 * Q*Q/r^2 = 9*10^9 * Q*Q / 0.02*0.02 = net

force on each = m * a
m a = 9*10^9 * Q*Q / 0.02*0.02
a = 225 m/s^2, m = 1*10^-3 kg

Q^2 = 0.02*0.02*225 * 10^-3 / 9*10^9
Q^2 = 10000 *10^-18
Q = 100 *10^-9
Q = 100 nC charge on each sphere <-----ANSWER